# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ Appendix F Proof of Proposition 1
Proof:
Problem (35) is equivalent to
$\begin{array}[]{cl}\displaystyle\max_{t,\bm{R}_{x},\bm{V},\bm{U}}&t\\ \text{s.t.}&\operatorname{Tr}\big{[}-\bm{V}+\bm{R}_{s}\big{]}\geq t\\ &\bm{V}\succeq\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}\\ &\operatorname{Tr}\left[{\bm{R}_{x}+\hat{\bm{R}}_{x}-2\bm{U}}\right]\leq\epsilon^{2}_{1}\\ &\hat{\bm{R}}^{1/2}_{x}\bm{R}_{x}\hat{\bm{R}}^{1/2}_{x}\succeq\bm{U}^{2}\\ &t\geq 0\\ &\bm{R}_{x}\succeq 0,~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x}\\ &\bm{V}\succeq\bm{0},~{}\bm{V}=\bm{V}^{\mathsf{H}}\\ &\bm{U}\succeq\bm{0},~{}\bm{U}=\bm{U}^{\mathsf{H}}.\end{array}$
Then, by applying the Schur complement, we have (36), which completes the proof. $\square$
Appendix G Proof of Proposition 3
@@
Note that $\bm{K}$ is invertible. The rest statements are due to Lemma 1. $\square$
Appendix H Additional Experimental Results on Scenario 2
In Section VI (in particular, Table I of Experiments) of the main body of the paper, we present the experimental results under the pilot data size of $20$. In this appendix, we display the results under the pilot data size of $10$.
TABLE II: Results of Scenario 2 (Pilot Size = 10)
Beamformer MSE Time Beamformer MSE Time
Wnr 80.59 6.96e-04 Wnr-DL 77.16 3.31e-05
Wnr-DR 51.35 3.67e+00 Wnr-CE 80.59 4.40e-04
Wnr-CE-DL 77.16 6.37e-05 Wnr-CE-DR 80.72 4.84e-04
Capon 98.31 4.05e-04 Capon-DL 95.56 6.91e-05
ZF 59.11 3.41e-04 Kernel 42.42 8.41e-04
Kernel-DL 40.85 9.20e-05
As we can see from Table I and Table II, a larger number of pilot data benefits the estimation performances of wireless signals. In addition, the diagonal loading operation can significantly improve the estimation performances especially when the pilot data size is relatively small. Moreover, nonlinear beamformers are preferable over linear ones when impulse channel noises exist.

### v2

@@ Appendix F Proof of Proposition 3
The rest claims are due to Lemma 1; NB: $\bm{K}$ is invertible. $\square$
Appendix G Additional Experimental Results
Complementary to experimental setups in Section VI, we consider pure complex Gaussian channel noises. First, we suppose that the transmit antennas emit continuous-valued complex signals; without loss of generality, Gaussian signals are used in experiments. The performance evaluation measure is therefore the mean-squared error (MSE). The experimental results are shown in Fig. 2.
(a) $N$=8, SNR 10dB, $\bm{R}_{v}$ Estimated
(b) $N$=8, SNR 10dB, $\bm{R}_{v}$ Known
(c) $N$=16, SNR 10dB, $\bm{R}_{v}$ Estimated
(d) $N$=16, SNR -10dB, $\bm{R}_{v}$ Estimated
Figure 2: Testing MSE against training pilot sizes under different numbers of receive antennas; only non-robust beamformers including non-diagonal-loading ones are considered. The true value of $\bm{R}_{v}$ can be unknown and estimated using pilot data. The signal-to-noise ratio (SNR) is $10$dB or $-10$dB.
From Fig. 2, the following main points can be outlined.
  1. 1.
    For a fixed number $M$ of transmit antennas, the larger the number $N$ of receive antennas, the smaller the MSE; cf. Figs. 2(a) and 2(c). This fact is well-established and is due to the benefit of antenna diversity. In addition, for fixed $N$ and $M$, the higher the SNR, the smaller the MSE; cf. Figs. 2(c) and 2(d); this is also well believed.
  2. 2.
    As the pilot size increases, the Wiener beamformer tends to have the best performance because the Wiener beamformer is optimal for the linear Gaussian signal model. When $\bm{R}_{v}$ is accurately known, the Wiener-CE beamformer outperforms the general Wiener beamformer (cf. Fig. 2(b)) because the former also exploits the information of the linear signal model in addition to the pilot data, while the latter only utilizes the pilot data. However, when $\bm{R}_{v}$ is estimated using the pilot data, the performances of the general Wiener beamformer and the Wiener-CE beamformer have no significant difference; cf. Figs. 2(a) and 2(c). Therefore, Fig. 2 validates our claim that channel estimation is not a necessary operation in receive beamforming and estimation of wireless signals; recall Subsection III-A3.
  3. 3.
    The ZF beamformer tends to be more efficient as $N$ increases; cf. Figs. 2(a) and 2(c). However, the ZF beamformer becomes less satisfactory when the SNR decreases; cf. Figs. 2(c) and 2(d). The Capon beamformer is also unsatisfactory when $N$ is small or the SNR is low.
  4. 4.
Second, we suppose that the transmit antennas emit discrete-valued symbols from a constellation that is modulated using quadrature phase-shift keying (QPSK). The performance evaluation measure is therefore the symbol error rate (SER). The experimental results are shown in Fig. 3. We find that all the conclusive main points from Fig. 2 can be obtained from Fig. 3 as well: this validates that minimizing MSE reduces SER. In addition, Figs. 3(c) and 3(d) reveal that the Wiener beamformer even slightly works better than the Wiener-CE beamformer when the pilot size is smaller than $15$ because the uncertainty in the estimated $\hat{\bm{R}}_{v}$, on the contrary, misleads the latter. Nevertheless, as the pilot size increases, the Wiener-CE beamformer tends to overlap the Wiener beamformer quickly.
(a) $N$=8, SNR 10dB, $\bm{R}_{v}$ Estimated
(b) $N$=8, SNR 10dB, $\bm{R}_{v}$ Known
(c) $N$=16, SNR 10dB, $\bm{R}_{v}$ Estimated
(d) $N$=16, SNR -10dB, $\bm{R}_{v}$ Estimated
Figure 3: Testing SER against training pilot sizes under different numbers of receive antennas; only non-robust beamformers including non-diagonal-loading ones are considered. The true value of $\bm{R}_{v}$ can be unknown and estimated using pilot data. The signal-to-noise ratio (SNR) is $10$dB or $-10$dB.